OpenStudy (adamk):

Verify the trig identity (1/2)sin4x = 2sinxcosx - 4(sin^3)xcosx No cross multiply. Must use trig identities.

OpenStudy (adamk):

I've got: 1/2sin4x = 2sinxcosx(1-2sin^2x) 1/2sin4x = sin2x(cos^2x)

OpenStudy (vishweshshrimali5):

Okay your first step was perfect... in second step check again that cos^2 x term... you did a mistake there

OpenStudy (vishweshshrimali5):

Recall that \(\sin^2 x= 1 - \cos^2 x\) .. make sure you didn't make any calculation mistake

OpenStudy (adamk):

Ah right. That should be cos2x.

OpenStudy (vishweshshrimali5):

Great! So we have... sin (2x) cos (2x)

OpenStudy (vishweshshrimali5):

Multiply and divide by 2

OpenStudy (adamk):

I don't understand that step.

OpenStudy (vishweshshrimali5):

Okay so we have.. sin(2x)cos(2x) I multiply with 2 and divide by 2 ... 1/2 * 2 * sin(2x) * cos(2x) = 1/2 * [2* sin(2x)*cos(2x)] The term in square brackets seem familiar?

OpenStudy (adamk):

Ah, OK that makes sense. Should it be 2 * (sin(2x)*cos(2x)) though?

OpenStudy (vishweshshrimali5):

Yeah.. 2 * sin(2x) * cos(2x) you can write as sin(4x)